Cost of Materials Used$60,000

Direct Labor costs $55,000

Factory Overhead $32,000

Work in Process, beg. $28,000

Work in Process, end. $18,000

Finished Goods,beg. $28,000

Finished Goods, end. $18,000

What is Cost of Goods Sold?
Answer #1

Total manufacturing cost=Direct materials used+Direct labor+Factory overhead

=(60,000+55,000+32,000)=147000

Cost of goods manufactured=Total manufacturing cost+Beginning WIP-Ending WIP

=147000+28,000-18,000=157000

Cost of goods sold=Cost of goods manufactured+Beginning Finished Goods-Ending Finished Goods

=157000+28,000-18,000

=$167000